Could you please help me with this?

A father leaves all of his money to his children when he dies. Child A
received 1/6 of the money. Child B receives 1/7 of the money. Child C receives
$23,490.00. How much money did Child A receive?

Answer by usyim88hk(158)

Frist we have to know the percentage of the money Child C gets
So it will be 1-(1/6)-(1/7) becuase child A and child B gets theirs and so the percentage left is for Child C
1-(1/6)-(1/7)
=(42/42)-(7/42)-(6/42)
=(42-7-6)/42
=(29/42)
So this is the percentage of the money Child C gets
and now we can divide 23490 by (42/29) to get the original amount of money that their dads left
23490/(42/29) = 34020
And we know that Child A recieved 1/6 of the money so now we can multiply 34020 by 1/6
34020/6 = 5670
So the total amount of money Child A recieved is $5670